OpenStudy (anonymous):

f(x)=x-1/x+2, g(x)=1/x

OpenStudy (ivettef365):

what is the question

OpenStudy (anonymous):

find function of f & g

OpenStudy (ivettef365):

what kind of function? could be (fog) of f(x)/g(x) or f(x) + g(x) needs to be more specific

OpenStudy (anonymous):

(fog) (x)= f(g(x))

OpenStudy (ivettef365):

ok, then all you need to do is substitute the x from the function f(x) with 1/x

OpenStudy (ivettef365):

you have f(x) = x-1/x-2 so it would be (1/x)-1/(1/x)-2

OpenStudy (anonymous):

would the answer be -x+1/x+2?

OpenStudy (ivettef365):

do you have choices

OpenStudy (anonymous):

im not sure if i did it right

OpenStudy (ivettef365):

the answer I get is the one that I showed you, don't think you can do anything else

OpenStudy (anonymous):

so it will be f(x)= (1/x)-1/(1/x)-2

OpenStudy (anonymous):

or i dont need to put f(x)?

OpenStudy (ivettef365):

actually now would be as follows: f(g(x)) = then the answer

OpenStudy (anonymous):

one last question about f(x)=x-1/x+2, g(x)=1/x, how come your answer change the sign to (1/x)-1/(1/x)-2, the +2, became -2

OpenStudy (ivettef365):

sorry you are right it is +2, it doesn't change, the only thing I did was substitute the x for 1/x

OpenStudy (anonymous):

thank you, got it! ur a big help

OpenStudy (ivettef365):

yw :)